OpenStudy (anonymous):

9x-3 > or equal to 51. explain.

OpenStudy (anonymous):

Explain what exactly?

OpenStudy (anonymous):

how to solve it.

OpenStudy (anonymous):

ok so what do u have to do to get 9x by itself?

OpenStudy (anonymous):

subtract 3 from 51 and -3

OpenStudy (alexandervonhumboldt2):

9x>=51+3 x>=54/9>=6

OpenStudy (alexandervonhumboldt2):

no add not subtarcrt

OpenStudy (anonymous):

oh right i thought you subtracted hold on

OpenStudy (anonymous):

add 3 then divide both sides by 9

Nnesha (nnesha):

@whysignupfor1question good job really appreciate it

OpenStudy (anonymous):

thanks and np

Nnesha (nnesha):

:)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

how do i divide both sides by 9, when i get 9x>54?

OpenStudy (anonymous):

oh it would be 6>9?

OpenStudy (anonymous):

9x/9=x 54/9=6

OpenStudy (anonymous):

x>6

OpenStudy (anonymous):

okay i have more questions ill tag you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nvm they are all graphs

OpenStudy (anonymous):

ok lol